OpenStudy (anonymous):

If I have a problem like this how do I solve it. x - 2/3= -5/6

OpenStudy (andriod09):

the problem is \[x-\frac {2} {3} = \frac {-5} {6}\]

OpenStudy (raden_zaikaria):

3(x - 2/3)= 3(-5/6)

OpenStudy (raden_zaikaria):

3x-2=-15/6

OpenStudy (raden_zaikaria):

3x-2=-5/2

OpenStudy (raden_zaikaria):

3x=-5/2 + 2

OpenStudy (raden_zaikaria):

3x=-5/2 + 4/2

OpenStudy (raden_zaikaria):

3x=-1/2

OpenStudy (raden_zaikaria):

x=-1/6

OpenStudy (anonymous):

-1/6